Assume that the population of heights of female college students is approximately normally distributed with mean m of 67.26 inches and standard deviation s of 5.96 inches. A random sample of 78 heights is obtained. Show all work.
(A) Find the mean and standard error of the x` distribution
(B) Find P(x>67.50)


Answer by stanbon(75887)

Assume that the population of heights of female college students is approximately normally distributed with mean m of 67.26 inches and standard deviation s of 5.96 inches. A random sample of 78 heights is obtained. Show all work.
(A) Find the mean and standard error of the x distribution
mean of sample means = mean of population = 67.26
std of sample means = 5.98/sqrt(78)
----------------------------------------

(B) Find P(x>67.50)
I am assuming "x" refers to the population.
z(67.5) = (67.5-67.26)/5.98 = 0.0401
---
P(x > 67.5) = P(z > 0.0401) = 0.4840
================
Cheers,
Stan H.